# Tree Posets: Supersaturation, Enumeration, and Randomness

### v2

@@ 1 Introduction
The question we are interested in this paper is one of supersaturation, that is, how many copies of a poset are we guaranteed once we are above the threshold of containing one. The simplest poset to consider the supersaturation question for are $2$-chains. Erdős and Katona conjectured that a family with $\binom{n}{\lfloor n/2\rfloor}+t$ sets in $[n]$ must contain at least $t\cdot\lfloor\frac{n+1}{2}\rfloor$ many $2$-chains. This conjecture was confirmed by Kleitman [23], who in fact showed that for every $0\leq a\leq 2^{n}$, every family in $[n]$ with size $a$ contains at least as many $2$-chains as the so-called centralized family of size $a$, i.e. a family of $a$ sets whose cardinalities are as close to $n/2+1/4$ as possible. Kleitman [23] further conjectured that the same families should also minimize the number of $k$-chains for every $k$. Five decades later, Kleitman’s result was rediscovered by Das, Gan, and Sudakov [11] and independently by Dove, Griggs, Kang, and Sereni [14]. Both papers further confirmed Kleitman’s conjecture for every $k$ and $a$ belonging to a certain range above the sum of the $k-1$ largest binomial coefficients. Subsequently, Balogh and Wagner [6] proved Kleitman’s conjecture for all $k$ and $a\leq(1-\varepsilon)2^{n}$, provided that $n$ is sufficiently large with respect to $k$ and $\varepsilon$. Finally, in a remarkable paper, Samotij [39] resolved Kleitman’s conjecture in full. There have been further generalizations of the supersaturation problem for $2$-chains to more general hosts other than the Boolean lattice such as the collection of subspaces of $\mathbb{F}_{q}^{n}$ ordered by set inclusion [32] or $\{0,1,2,\dots r\}^{n}$ [6, 32]. The latter problem gives rise to a natural generalization of Kleitman’s problem, and while weaker approximate results do hold for $r\geq 2$ as shown by Noel, Sudakov and Scott [32], the exact analogue of Kleitman’s conjecture (see  [6, 32]) fails as shown by Balogh, Petříčková, and Wagner [7].
@@
Let P, Q be two finite posets, that is, they are finite sets equipped with partial orders $<_{P}$ and $<_{Q}$. A poset homomorphism is a function $f:P\to Q$ such that $f(x)<_{Q}f(y)$ whenever $x<_{P}y$. An induced poset homomorphism is a function $f:P\to Q$ such that $f(x)<_{Q}f(y)$ if and only if $x<_{P}y$. We say that a poset $Q$ contains another poset $P$ if there is an injective poset homomorphism from $P$ to $Q$. We say that a poset $Q$ contains an induced copy of another poset $P$ if there is an injective induced poset homomorphism from $P$ to $Q$. If a poset $Q$ does not contain a copy of another poset $P$, we say that $Q$ is $P$-free. If $Q$ does not contain an induced copy of $P$, we say that $Q$ is induced $P$-free. Given a poset $P$ and an integer $n$, we define $La(n,P)$ to be the largest size of a $P$-free subfamily of $\mathcal{B}_{n}$ and $La^{*}(n,P)$ the largest size of an induced $P$-free subfamily of $\mathcal{B}_{n}$. Motivated by a number of early results in the study of $La(n,P)$ (see for example [12, 13, 19, 42]), Griggs and Lu [21] and independently Bukh [9] made the following conjecture on the form of $La(n,P)$.
@@
In the setting of posets, containers have been most widely used in counting subsets of $\mathcal{B}_{n}$ avoiding $2$-chains, which is equivalent to the problem of counting antichains. When the host is the Boolean lattice, this problem was solved independently by Balogh, Treglown, and Wagner [5] and by Noel, Sudakov, and Scott [32]. The second group [32] also generalized this to hosts being subspaces of $\mathbb{F}_{n}^{q}$ ordered by inclusion, and sets of divisors of a square of a square free integer. Note that the latter is equivalent to studying $\{0,1,2\}^{n}$ under the natural ordering. In the more general setting of counting antichains in $[t]^{n}$, which turns out to be connected to a Ramsey-theoretic question in ordered hypergraphs [31], there has been much recent progress, most notably by Pohoata and Zakharov [37], Park, Sarantis, and Tetali [34] and finally, by Falgrav-Ravry, Rävy, and Tomon [17].
@@
Finally, it is worth highlighting that the main tool of this paper Theorem 4.15 which we believe is of independent interest, and can be viewed as a general tool similar to results in graphs of passing to subgraphs of high minimum degree. While the exact analogue of such a result is out of reach, that is, obtaining a subfamily in the poset setting of “high minimum degree” we obtain a sequence of nested subfamilies of our original family $\mathcal{F}$ having high minimum degree in the predecessor of the sequence which still allows us to greedily find many embeddings of a tree poset, similar to embedding copies of a fixed tree in a high minimum degree subgraph. We believe that this result will have further applications in the future.
@@ 2 Notation
~$\widetilde{\mathcal{B}}_{n}:=\left\{F\in\mathcal{B}_{n}:\left|F-\frac{n}{2}\right|<2\sqrt{n\ln n}\right\}.$
Using some standard tools, we will assume that the family $\mathcal{F}$ we are working with is subfamily of ~$\widetilde{\mathcal{B}}_{n}$ instead of $\mathcal{B}_{n}$.
@@
$\mathcal{L}^{i}(\mathcal{M})=\{D\in\mathcal{B}_{n}:D\text{ is the }i\text{th member of a }q\text{-chain in }\mathcal{M}\}.$
Our proof will rely on obtaining a nice nested sequence of $q$-marked chains from $\mathcal{F}$ that are of a specific form. To this end, given a family $\mathcal{T}$ of $1$-marked chains of some subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ and $\chi\in\mathcal{C}$, we define $\mathcal{T}(\chi)=\{F:(\chi,F)\in\mathcal{T}\}$. We say that $\mathcal{T}$ is $q$-strong if for each $\chi\in\mathcal{C}$ where $\mathcal{T}(\chi)\neq\emptyset$, we have that $|\mathcal{T}(\chi)|\geq q$. For a $q$-strong $1$-marked chain family $\mathcal{T}$ from $\mathcal{F}$, we define the $q$-th power of $\mathcal{T}$, denoted by $\mathcal{T}[q]$, to be the $q$-marked chain family
@@
Using Chernoff bounds, it is standard [8, 21] to show that the number of sets $F\in\mathcal{B}_{n}$ with $||F|-n/2|>2\sqrt{n\ln{n}}$ is $o({n\choose n/2})$. Thus, whenever we are given a dense subfamily $\mathcal{F}$ of $\mathcal{B}_{n}$ i.e. $|\mathcal{F}|=\Omega({n\choose n/2})$, by leaving out at most $o(\binom{n}{n/2})$ members of $\mathcal{F}$, we may assume that all $F\in\mathcal{F}$ lie in the family ~$\widetilde{\mathcal{B}}_{n}$.
At the heart of all of our arguments is a general result Theorem 4.15 which says that any subset ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ of large Lubell weight, more precisely at least of weight at least $q-1+\varepsilon$, contains a nested sequence of large $q$-marked chain families that have a certain robustness property which we will describe momentarily. Note that if $\mathcal{F}$ is of size at least $(q-1+\varepsilon){n\choose\lfloor n/2\rfloor}$ then the hypothesis above on the Lubell weight is satisfied.
@@ 4.1 Counting Lemmas on Marked Chains
Then, $\mathcal{T}$ is a $q$-strong $1$-marked chain family from $\mathcal{F}$, satisfying $|\mathcal{T}|\geq\varepsilon n!$. In particular, if $|\mathcal{F}|\geq(q-1+\varepsilon)\max_{F\in\mathcal{F}}{n\choose|F|}$ then $|\mathcal{T}|\geq\varepsilon n!$
@@
Let =$\mathcal{M}=\{(\chi,F):\chi\in\mathcal{C},F\in\mathcal{F}\}$. For each $i\in[n]$, let $C_{i}$ denote the number of full chains $\chi$ that contain exactly $i$ members of $\mathcal{F}$. Then $|\mathcal{M}|=\sum_{i=1}^{n}iC_{i}$. On the other hand, for each $F\in\mathcal{F}$, the number of full chains in $\mathcal{B}_{n}$ that contain $F$ is exactly $\frac{n!}{\binom{n}{|F|}}$. Hence,
@@
Let $\mathcal{F}$ be a subfamily of $\mathcal{B}_{n}$ and $q$ a positive integer. Let $\mathcal{T}$ be a $q$-strong $1$-marked chain family from $\mathcal{F}$ and let $\mathcal{M}=\mathcal{T}[q]$ be the $q$-th power of $\mathcal{T}$. For each $i\in[q]$ and each $\chi\in\mathcal{C}$ let
$\mathcal{L}^{i}(\mathcal{M},\chi):=\{F\in\mathcal{F}:\exists(\chi,Q)\in\mathcal{M}\text{ such that $F$ is the $i$-th member on $Q$}\},$
@@
Consider any $\chi\in\mathcal{C}$ where $\mathcal{T}(\chi)\neq\emptyset$. Because $\mathcal{T}$ is $q$-strong, we have $|\mathcal{T}(\chi)|\geq q$ by definition. Note that for each $F\in\mathcal{T}(\chi)$ that is not among the largest $i-1$ members or smallest $q-i-1$ members in $\mathcal{T}(\chi)$, there exists a $q$-chain in $\binom{\mathcal{T}(\chi)}{q}$ that contains $F$ as the $i$-th member. Hence, $|\mathcal{L}^{i}(\mathcal{M},\chi)|\geq\frac{1}{q}|\mathcal{T}(\chi)|$ and therefore
@@
On the other hand, for each $F\in\mathcal{L}^{i}(\mathcal{M})$, $F$ is contained in exactly $n!/\binom{n}{|F|}$ full chains of $\mathcal{B}_{n}$ and hence belongs to $\mathcal{L}^{i}(\mathcal{M},\chi)$ for at most $n!/\binom{n}{|F|}$ different $\chi$. Hence,
@@ 3. $(\hat{c})$
For all integers $q\geq 1$ and for all reals $\varepsilon>0$, there exists some $\delta>0$ such that the following holds. Let ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ and let $\mathcal{T}^{0}$ be a $q$-strong $1$-marked chain family with markers from $\mathcal{F}$ such that $|\mathcal{T}^{0}|\geq\varepsilon n!$. Then there exists a collection of $1$-marked chains $\mathcal{T}^{|P|}\subseteq\mathcal{T}^{|P|-1}\subseteq\dots\subseteq\mathcal{T}^{0}$ satisfying the following:
@@
    For each $j=1,\dots,|P|$, for each $(\chi,Q)\in\mathcal{T}^{j}[q]$ and $F\in Q$, $F$ is $\delta$-robust with respect to $\mathcal{T}^{j-1}[q]$.
@@
Before starting the proof, we set up some notation and claims that will be useful. In what follows, we fix ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ as in the theorem statement together with the large constant
@@
We begin by explicitly defining $\vec{b}(\chi)$ for each $\chi\in\mathcal{C}_{1}^{j}(\downarrow)$, and will refer to this vector as the lower-bad profile for $\chi$ relative to $\mathcal{M}^{j-1}$
@@
In particular, this implies there are at least $q-i$ additional members of $\mathcal{T}^{j-1}(\chi)$ on $\chi$ below $F_{a_{d}}$. Because $F_{a_{d}}\in B^{j-1}(\chi,i,\downarrow)$, it is the $i$th member of some $q$-marked chain in $\mathcal{M}^{j-1}(\chi)=\binom{\mathcal{T}^{j-1}(\chi)}{q}$, and in particular there exists at least $i-1$ members which come before it, and at least $q-i$ members coming after in inside $\mathcal{T}^{j-1}(\chi)$. By property $(b)$ of witness sets we have then $\{F_{a_{d+1}},\dots F_{a_{d+q-i}}\}\cap\mathcal{W}(F_{a_{d}},i,\mathcal{T}^{(j-1)}[q])\neq\emptyset$. Let $b_{1}=a_{d}$ and let $b_{1}^{\prime}$ be the index of the any member $\{F_{a_{d+1}},\dots F_{a_{d+q-i}}\}\cap\mathcal{W}(F_{a_{d}},i,\mathcal{T}^{(j-1)}[q])$.
Now let $d^{\prime}$ be the smallest integer such that $F_{a_{d^{\prime}}}\in B^{j-1}(\chi,i,\downarrow)$ and $a_{d^{\prime}}>b_{1}^{\prime}$ if it exists. Just as before, we are guaranteed one of $F_{a_{d^{\prime}+1}},\dots F_{a_{d^{\prime}+q-i}}$ belongs to $\mathcal{W}(F_{a_{d^{\prime}}},i,\mathcal{T}^{(j-1)}[q])$, and we let $b_{2}=a_{d^{\prime}}$ and $b_{2}^{\prime}$ be the index of the any member of $\mathcal{W}(F_{a_{d^{\prime}}},i,\mathcal{T}^{(j-1)}[q])\cap\{F_{a_{d^{\prime}+1}},\dots F_{a_{d^{\prime}+q-i}}\}$. We continue to repeat the process, e.g. by defining $d^{\prime\prime}$ to be the smallest integer such that $F_{a_{d^{\prime\prime}}}\in B^{j-1}(\chi,i,\downarrow)$ and $a_{d^{\prime\prime}}>b_{2}^{\prime}$, if it exists, until no more choices remain. Note that the process goes on at least $\frac{1}{q-i}|B^{j-1}(\chi,i,\downarrow)|$ many steps, as between $F_{b_{i}}$ and $F_{b_{i}^{\prime}}$ there are at most $q-i$ members of $B^{j-1}(\chi,i,\downarrow)$. Furthermore, $\vec{b}(\chi)\subseteq[|\mathcal{T}^{0}(\chi)|]$.
@@
By our assumption of $\chi\in{\mathcal{C}}_{1}^{j}(\downarrow)$, we must have $\chi\in{\mathcal{C}}^{\ell}_{2}$ for $\ell=0,1,\dots,j-1$ and $\chi\in{\mathcal{C}}_{1}^{j}(\downarrow)$. Therefore, for $11\leq\ell\leq j-$ 1, we know that $|\mathcal{T}^{\ell}(\chi)|\geq(1-\frac{2}{\Delta})|\mathcal{T}^{\ell-1}(\chi)|$, and hence $|\mathcal{T}^{0}(\chi)|\leq(\frac{\Delta}{\Delta-2})^{j}|\mathcal{T}^{j-1}(\chi)|$. By definition of $\chi\in\mathcal{C}_{1}^{j}(\downarrow)$, we know that $|B^{j-1}(\chi,\downarrow)|\geq|{\mathcal{T}}^{j-1}(\chi)|/\Delta$. By our choice of $i$, $|B^{j-1}(\chi,i,\downarrow)|\geq\frac{1}{q}|B^{j-1}(\chi,\downarrow)|$. By our earlier observation, we have that $m\geq\frac{1}{q-i}|B^{j-1}(\chi,i,\downarrow)|$. It therefore follows that
@@
For all integers $q\geq 1$ and reals $\varepsilon>0$, there exists some $\delta>0$ such that the following holds. If ~$\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ is a family with $\mu(\mathcal{F})\geq q-1+\varepsilon$, then there exists a nested sequence of $q$-marked chains $\mathcal{M}^{|P|}\subseteq\mathcal{M}^{|P|-1}\subseteq\cdots\subseteq\mathcal{M}^{0}$ such that for all $j\in[|P|]$ we have:

### v3

@@ Abstract
Keywords: tree poset, chain, container lemma, enumeration, supersaturation.
2020 MSC subject classification: 05D05.
@@ 1 Introduction
The question we are interested in this paper is one of supersaturation, that is, how many copies of a poset are we guaranteed once we are above the threshold of containing one. The simplest poset to consider the supersaturation question for are $2$-chains. Erdős and Katona conjectured that a family with $\binom{n}{\lfloor n/2\rfloor}+t$ sets in $[n]$ must contain at least $t\cdot\lfloor\frac{n+1}{2}\rfloor$ many $2$-chains. This conjecture was confirmed by Kleitman [23], who in fact showed that for every $0\leq a\leq 2^{n}$, every family in $[n]$ with size $a$ contains at least as many $2$-chains as the so-called centralized family of size $a$, i.e. a family of $a$ sets whose cardinalities are as close to $n/2+1/4$ as possible. Kleitman [23] further conjectured that the same families should also minimize the number of $k$-chains for every $k$. Five decades later, Kleitman’s result was rediscovered by Das, Gan, and Sudakov [11] and independently by Dove, Griggs, Kang, and Sereni [14]. Both papers further confirmed Kleitman’s conjecture for every $k$ and $a$ belonging to a certain range above the sum of the k1k-1 largest binomial coefficients. Subsequently, Balogh and Wagner [6] proved Kleitman’s conjecture for all $k$ and $a\leq(1-\varepsilon)2^{n}$, provided that $n$ is sufficiently large with respect to $k$ and $\varepsilon$. Finally, in a remarkable paper, Samotij [39] resolved Kleitman’s conjecture in full. There have been further generalizations of the supersaturation problem for $2$-chains to more general hosts other than the Boolean lattice such as the collection of subspaces of $\mathbb{F}_{q}^{n}$ ordered by set inclusion [32] or $\{0,1,2,\dots r\}^{n}$ [6, 32]. The latter problem gives rise to a natural generalization of Kleitman’s problem, and while weaker approximate results do hold for $r\geq 2$ as shown by Noel, Sudakov and Scott [32], the exact analogue of Kleitman’s conjecture (see  [6, 32]) fails as shown by Balogh, Petříčková, and Wagner [7].
@@
Let P, Q be two finite posets, that is, they are finite sets equipped with partial orders $_{P}$ and $_{Q}$. A poset homomorphism is a function $f:P\to Q$ such that $f(x)<_{Q}f(y)$ whenever $x<_{P}y$. An induced poset homomorphism is a function $f:P\to Q$ such that $f(x)<_{Q}f(y)$ if and only if $x<_{P}y$. We say that a poset $Q$ contains another poset $P$ if there is an injective poset homomorphism from $P$ to $Q$. We say that a poset $Q$ contains an induced copy of another poset $P$ if there is an injective induced poset homomorphism from $P$ to $Q$. If a poset $Q$ does not contain a copy of another poset $P$, we say that $Q$ is $P$-free. If $Q$ does not contain an induced copy of $P$, we say that $Q$ is induced $P$-free. Given a poset $P$ and an integer $n$, we define $La(n,P)$ to be the largest size of a $P$-free subfamily of $\mathcal{B}_{n}$ and $La^{*}(n,P)$ the largest size of an induced $P$-free subfamily of $\mathcal{B}_{n}$. Motivated by a number of early results in the study of $La(n,P)$ (see for example [12, 13, 19, 42]), Griggs and Lu [21] and independently Bukh [9] made the following conjecture on the form of $La(n,P)$.
@@
In the setting of posets, containers have been most widely used in counting subsets of $\mathcal{B}_{n}$ avoiding $2$-chains, which is equivalent to the problem of counting antichains. When the host is the Boolean lattice, this problem was solved independently by Balogh, Treglown, and Wagner [5] and by Noel, Sudakov, and Scott [32]. The second group [32] also generalized this to hosts being subspaces of $\mathbb{F}_{n}^{q}$ ordered by inclusion, and sets of divisors of a square of a square free integer. Note that the latter is equivalent to studying $\{0,1,2\}^{n}$ under the natural ordering. In the more general setting of counting antichains in $t]^{n}$, which turns out to be connected to a Ramsey-theoretic question in ordered hypergraphs [31], there has been much recent progress, most notably by Pohoata and Zakharov [37], Park, Sarantis, and Tetali [34] and finally, by Falgrav-Ravry, Rävy, and Tomon [17].
@@
Finally, it is worth highlighting that the main tool of this paper, Theorem 4.15, which we believe is of independent interest, and can be viewed as a general tool similar to results in graphs of passing to subgraphs of high minimum degree. While the exact analogue of such a result is out of reach, that is, obtaining a subfamily in the poset setting of “high minimum degree” we obtain a sequence of nested subfamilies of our original family $\mathcal{F}$ having high minimum degree in the predecessor of the sequence which still allows us to greedily find many embeddings of a tree poset, similar to embedding copies of a fixed tree in a high minimum degree subgraph. We believe that this result will have further applications in the future.
@@ 2 Notation
$\widetilde{\mathcal{B}}_{n}:=\left\{F\in\mathcal{B}_{n}:\left|F-\frac{n}{2}\right|<2\sqrt{n\ln n}\right\}.$
Using some standard tools, we will assume that the family $\mathcal{F}$ we are working with is subfamily of $\widetilde{\mathcal{B}}_{n}$ instead of $\mathcal{B}_{n}$.
@@
[…]
Our proof will rely on obtaining a nice nested sequence of $q$-marked chains from $\mathcal{F}$ that are of a specific form. To this end, given a family $\mathcal{T}$ of $1$-marked chains of some subfamily $\mathcal{F}\subseteq\mathcal{B}_{n}$ and $\chi\in\mathcal{C}$, we define $\mathcal{T}(\chi)=\{F:(\chi,F)\in\mathcal{T}\}$. We say that $\mathcal{T}$ is $q$-strong if for each $\chi\in\mathcal{C}$ where $\mathcal{T}(\chi)\neq\emptyset$, we have that $|\mathcal{T}(\chi)|\geq q$. For a $q$-strong 11-marked chain family $\mathcal{T}$ from $\mathcal{F}$, we define the $q$-th power of $\mathcal{T}$, denoted by $\mathcal{T}[q]$, to be the $q$-marked chain family
@@
Using Chernoff bounds, it is standard [8, 21] to show that the number of sets $F\in\mathcal{B}_{n}$ with $||F|-n/2|>2\sqrt{n\ln{n}}$ is $o({n\choose n/2})$. Thus, whenever we are given a dense subfamily $\mathcal{F}$ of $\mathcal{B}_{n}$ i.e. $|\mathcal{F}|=\Omega({n\choose n/2})$, by leaving out at most $o(\binom{n}{n/2})$ members of $\mathcal{F}$, we may assume that all $F\in\mathcal{F}$ lie in the family $\widetilde{\mathcal{B}}_{n}$.
At the heart of all of our arguments is a general result Theorem 4.15 which says that any subset $\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ of large Lubell weight, more precisely at least of weight at least $q-1+\varepsilon$, contains a nested sequence of large $q$-marked chain families that have a certain robustness property which we will describe momentarily. Note that if $\mathcal{F}$ is of size at least $(q-1+\varepsilon){n\choose\lfloor n/2\rfloor}$, then the hypothesis above on the Lubell weight is satisfied.
@@ 4.1 Counting Lemmas on Marked Chains
Then, $\mathcal{T}$ is a $q$-strong 11-marked chain family from $\mathcal{F}$, satisfying $|\mathcal{T}|\geq\varepsilon n!$. In particular, if $|\mathcal{F}|\geq(q-1+\varepsilon)\max_{F\in\mathcal{F}}{n\choose|F|}$ then $|\mathcal{T}|\geq\varepsilon n!$
@@
Let $\mathcal{M}=\{(\chi,F):\chi\in\mathcal{C},F\in\mathcal{F}\}$. For each $i\in[n]$, let $C_{i}$ denote the number of full chains $\chi$ that contain exactly $i$ members of $\mathcal{F}$. Then $|\mathcal{M}|=\sum_{i=1}^{n}iC_{i}$. On the other hand, for each $F\in\mathcal{F}$, the number of full chains in $\mathcal{B}_{n}$ that contain $F$ is exactly $\frac{n!}{\binom{n}{|F|}}$. Hence,
@@
Let $\mathcal{F}$ be a subfamily of $\mathcal{B}_{n}$ and $q$ a positive integer. Let $\mathcal{T}$ be a $q$-strong 11-marked chain family from $\mathcal{F}$ and let $\mathcal{M}=\mathcal{T}[q]$ be the $q$-th power of $\mathcal{T}$. For each $i\in[q]$ and each $\chi\in\mathcal{C}$ let
[…]
@@
Consider any $\chi\in\mathcal{C}$ where $\mathcal{T}(\chi)\neq\emptyset$. Because $\mathcal{T}$ is $q$-strong, we have $|\mathcal{T}(\chi)|\geq q$ by definition. Note that for each $F\in\mathcal{T}(\chi)$ that is not among the largest i1i-1 members or smallest qi1q-i-1 members in $\mathcal{T}(\chi)$, there exists a $q$-chain in $\binom{\mathcal{T}(\chi)}{q}$ that contains $F$ as the $i$-th member. Hence, $|\mathcal{L}^{i}(\mathcal{M},\chi)|\geq\frac{1}{q}|\mathcal{T}(\chi)|$ and therefore
@@
On the other hand, for each $\in\mathcal{L}^{i}(\mathcal{M}), FF$ is contained in exactly $n!/\binom{n}{|F|}$ full chains of $\mathcal{B}_{n}$ and hence belongs to $\mathcal{L}^{i}(\mathcal{M},\chi)$ for at most $n!/\binom{n}{|F|}$ different $\chi$. Hence,
@@ 3. $(c)$
For all integers $q\geq 1$ and for all reals $\varepsilon>0$, there exists some $\delta>0$ such that the following holds. Let $\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ and let $\mathcal{T}^{0}$ be a $q$-strong 11-marked chain family with markers from $\mathcal{F}$ such that $|\mathcal{T}^{0}|\geq\varepsilon n!$. Then there exists a collection of $1$-marked chains $\mathcal{T}^{|P|}\subseteq\mathcal{T}^{|P|-1}\subseteq\dots\subseteq\mathcal{T}^{0}$ satisfying the following:
@@
    For each $j=1,\dots,|P|$, for each $(\chi,Q)\in\mathcal{T}^{j}[q]$ and $\in Q, FF$ is $\delta$-robust with respect to $\mathcal{T}^{j-1}[q]$.
@@
Before starting the proof, we set up some notation and claims that will be useful. In what follows, we fix $\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ as in the theorem statement together with the large constant
@@
We begin by explicitly defining $\vec{b}(\chi)$ for each $\chi\in\mathcal{C}_{1}^{j}(\downarrow)$, and will refer to this vector as the lower-bad profile for $\chi$ relative to $\mathcal{M}^{j-1}$ .
@@
In particular, this implies there are at least qiq-i additional members of $\mathcal{T}^{j-1}(\chi)$ on $\chi$ below $F_{a_{d}}$. Because $F_{a_{d}}\in B^{j-1}(\chi,i,\downarrow)$, it is the $i$th member of some $q$-marked chain in $\mathcal{M}^{j-1}(\chi)=\binom{\mathcal{T}^{j-1}(\chi)}{q}$, and in particular there exists at least i1i-1 members which come before it, and at least qiq-i members coming after in inside $\mathcal{T}^{j-1}(\chi)$. By property $(b)$ of witness sets we have then $\{F_{a_{d+1}},\dots F_{a_{d+q-i}}\}\cap\mathcal{W}(F_{a_{d}},i,\mathcal{T}^{(j-1)}[q])\neq\emptyset$. Let $b_{1}=a_{d}$ and let $b_{1}^{\prime}$ be the index of the any member $\{F_{a_{d+1}},\dots F_{a_{d+q-i}}\}\cap\mathcal{W}(F_{a_{d}},i,\mathcal{T}^{(j-1)}[q])$.
Now let $d^{\prime}$ be the smallest integer such that $F_{a_{d^{\prime}}}\in B^{j-1}(\chi,i,\downarrow)$ and $a_{d^{\prime}}>b_{1}^{\prime}$ if it exists. Just as before, we are guaranteed one of $F_{a_{d^{\prime}+1}},\dots F_{a_{d^{\prime}+q-i}}$ belongs to $\mathcal{W}(F_{a_{d^{\prime}}},i,\mathcal{T}^{(j-1)}[q])$, and we let $b_{2}=a_{d^{\prime}}$ and $b_{2}^{\prime}$ be the index of the any member of $\mathcal{W}(F_{a_{d^{\prime}}},i,\mathcal{T}^{(j-1)}[q])\cap\{F_{a_{d^{\prime}+1}},\dots F_{a_{d^{\prime}+q-i}}\}$. We continue to repeat the process, e.g. by defining $d^{\prime\prime}$ to be the smallest integer such that $F_{a_{d^{\prime\prime}}}\in B^{j-1}(\chi,i,\downarrow)$ and $a_{d^{\prime\prime}}>b_{2}^{\prime}$, if it exists, until no more choices remain. Note that the process goes on at least $\frac{1}{q-i}|B^{j-1}(\chi,i,\downarrow)|$ many steps, as between $F_{b_{i}}$ and $F_{b_{i}^{\prime}}$ there are at most qiq-i members of $B^{j-1}(\chi,i,\downarrow)$. Furthermore, $\vec{b}(\chi)\subseteq[|\mathcal{T}^{0}(\chi)|]$.
@@
By our assumption of $\chi\in{\mathcal{C}}_{1}^{j}(\downarrow)$, we must have $\chi\in{\mathcal{C}}^{\ell}_{2}$ for $\ell=0,1,\dots,j-1$ and $\chi\in{\mathcal{C}}_{1}^{j}(\downarrow)$. Therefore, for $1\leq\ell\leq j-1$, we know that $|\mathcal{T}^{\ell}(\chi)|\geq(1-\frac{2}{\Delta})|\mathcal{T}^{\ell-1}(\chi)|$, and hence $|\mathcal{T}^{0}(\chi)|\leq(\frac{\Delta}{\Delta-2})^{j}|\mathcal{T}^{j-1}(\chi)|$. By definition of $\chi\in\mathcal{C}_{1}^{j}(\downarrow)$, we know that $|B^{j-1}(\chi,\downarrow)|\geq|{\mathcal{T}}^{j-1}(\chi)|/\Delta$. By our choice of $i$, $|B^{j-1}(\chi,i,\downarrow)|\geq\frac{1}{q}|B^{j-1}(\chi,\downarrow)|$. By our earlier observation, we have that $m\geq\frac{1}{q-i}|B^{j-1}(\chi,i,\downarrow)|$. It therefore follows that
@@
For all integers $q\geq 1$ and reals $\varepsilon>0$, there exists some $\delta>0$ such that the following holds. If $\mathcal{F}\subseteq\widetilde{\mathcal{B}}_{n}$ is a family with $\mu(\mathcal{F})\geq q-1+\varepsilon$, then there exists a nested sequence of $q$-marked chains $\mathcal{M}^{|P|}\subseteq\mathcal{M}^{|P|-1}\subseteq\cdots\subseteq\mathcal{M}^{0}$ such that for all $j\in[|P|]$ we have: